OpenStudy (anonymous):

x^2+x=0 solve by factoring?

hartnn (hartnn):

take x common x(x+1)=0 can u solve further?

OpenStudy (anonymous):

x=1 ?

hartnn (hartnn):

nopes, x(x+1) =0 x=0 OR x+1=0 so what are the two values of x ?

OpenStudy (anonymous):

0 or 1

hartnn (hartnn):

if x+1=0 then x will be -1, isn't it ?

OpenStudy (anonymous):

oh yeah! you right. thanks a bunch!

hartnn (hartnn):

welcome :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)